OpenStudy (idku):

another question power series rep. for tan^(-1)x

OpenStudy (misty1212):

differentiate get \[\frac{1}{1+x^2}\] find the power series, integrate term by term

OpenStudy (idku):

I know that \[\frac{d}{dx}\left(\tan^{-1}x\right)=\frac{1}{1+x^2}\] (this can be proven tan^(-1)x=y tan(y)=x derivative, dy/dx * sec^2(y) = 1 dy/dx * ( tan^2(y) + 1) = 1 dy/dx = 1/(1+tan^2y) dy/dx = 1/(1+x^2) the prove is here, that is no the concern) BUT, \[\frac{1}{1-(-x^2)}=\sum_{n=1}^{\infty}x^{2n}\]\[\tan^{-1}(x)=\sum_{n=1}^{\infty}\frac{x^{2n+1}}{2n+1}\]

OpenStudy (idku):

right ?

OpenStudy (misty1212):

you got this !

OpenStudy (idku):

oh, I forgot something

OpenStudy (idku):

\[\frac{1}{1-(-x^2)}=\sum_{n=1}^{\infty}(-1)^{n}x^{2n}\]

OpenStudy (idku):

\[\tan^{-1}x=\sum_{n=1}^{\infty}(-1)^{n}\frac{x^{2n+1}}{2n+1}\]

OpenStudy (idku):

like this... correct ?

OpenStudy (phi):

again, the lower limit is n=0 not n=1

OpenStudy (idku):

oh, so all the work is correct besides that it is n=0 (and not n=1) ?

OpenStudy (phi):

yes

OpenStudy (idku):

tnx